OpenStudy (bananas1234):

Which expression is equivalent to –(–2m + 1)? –3m 3m –2m – 1 2m – 1

OpenStudy (ahsome):

\[-(-2m+1)\]Is the same as: \[-1\times(-2m+1)\] \[(-2m\times-1)+(1\times-1)\]\[2m+-1\]\[2m-1\]

OpenStudy (bananas1234):

That is pretty simple! Thanks!

OpenStudy (ahsome):

No problem @Bananas1234 :D